OpenStudy (anonymous):

Evaluate logbase3 (3/y)

OpenStudy (turingtest):

\[\log(\frac ab)=\log a-\log b\]

OpenStudy (anonymous):

logbase3 3-logbase3 y?

OpenStudy (turingtest):

yep :) and what is\[\log_3(3)\]?

OpenStudy (anonymous):

1

OpenStudy (turingtest):

so this simplifies to...?

OpenStudy (anonymous):

1+logY?

OpenStudy (turingtest):

I think the rule is\[\log(\frac ab)=\log a-\log b\]careful with your signs!

OpenStudy (anonymous):

OR 1-LOGBASE3 Y?

OpenStudy (turingtest):

yep :)